OpenStudy (anonymous):

If sin x = 0.20, cos x = .98, sin y = .12, and cos y = .99, which of the following is cos(x + y)

OpenStudy (raden):

use the identity : cos(x + y) = cos(x)cos(y) - sin(x)sin(y)

OpenStudy (raden):

therefore, cos(x + y) = 0.98(0.99) - 0.20(0.12) = .... simplify that

OpenStudy (anonymous):

i didnt get an available answer @RadEn

OpenStudy (anonymous):

A. .07 B. .78 C. .70 D. .80

OpenStudy (raden):

maybe something typo one of the values given above ??